Effect of Financing on Earnings per Share

Domanico Co., which produces and sells biking equipment, is financed as follows:

Bonds payable, 10% (issued at face amount) $800,000
Preferred $2 stock, $20 par 800,000
Common stock, $25 par 800,000

Income tax is estimated at 40% of income.

Determine the earnings per share on common stock, assuming that the income before bond interest and income tax is (a) $320,000, (b) $400,000, and (c) $480,000.

Enter answers in dollars and cents, rounding to two decimal places.

a. Earnings per share on common stock $

b. Earnings per share on common stock $

c. Earnings per share on common stock $
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statement showing Computations
Particulars a) 320,000
EBIT            320,000.00
Less Interest on Bond = 800,000 * 10%              80,000.00
EBT            240,000.00
Less Income tax at 40%              96,000.00
EAT            144,000.00
Less Preferred Dividend = 800,000/20*2              80,000.00
Earning for common stock shareholders              64,000.00
Common Stock shares = 800,000/25              32,000.00
Earnings per share on common stock                         2.00
b)
Particulars b) 400,000
EBIT            400,000.00
Less Interest on Bond = 800,000 * 10%              80,000.00
EBT            320,000.00
Less Income tax at 40%            128,000.00
EAT            192,000.00
Less Preferred Dividend = 800,000/20*2              80,000.00
Earning for common stock shareholders            112,000.00
Common Stock shares = 800,000/25              32,000.00
Earnings per share on common stock                         3.50
c)
Particulars c) 480,000
EBIT            480,000.00
Less Interest on Bond = 800,000 * 10%              80,000.00
EBT            400,000.00
Less Income tax at 40%            160,000.00
EAT            240,000.00
Less Preferred Dividend = 800,000/20*2              80,000.00
Earning for common stock shareholders            160,000.00
Common Stock shares = 800,000/25              32,000.00
Earnings per share on common stock                         5.00